# 15.1cm25.0cm

### v2

@@ 1. Introduction
It is noteworthy that recombination can weaken the selective disadvantage of a pathogenic recessive allele by exporting mutations to other members of the population [7, 13, 15]. This is particularly effective for mutations with a low dominance coefficient $h$ ($h=0$ is usually referred to as recessivity) because the effect on the phenotypic level requires them to occur in a homozygous or compound heterozygous state. In fact, the genetic load, which is defined as the reduction in population fitness compared to a theoretical “perfectly adapted” genotype, also increases with effective population size [11, 17]. Therefore, Charlesworth, Kondrashov, and others hypothesized that recombination is a force that helps to balance genetic load to a certain degree in growing populations or genomes with an increasing number of recessive genes [4, 16].
@@
When recombination is absent, deleterious mutations accumulate, and the haplotypes without mutations are lost due to random genetic drift. Muller first described this stochastic process as a ratchet in which each click represents the irreversible loss of the least loaded class from the population [10, 21]. By theoretical arguments, Charlesworth and Charlesworth argued that in diploid genomes the accumulation of pathogenic or lethal recessive alleles can result in the “cristalization” of the population, that is, the occurrence of haplotypes that are incompatible with each other [5]. In our work, we study the stochastic process of Muller‘s ratchet in the limit of strong selection on completely recessive diploid genomes in a parameter space that is close to the barrier drift hypothesis. We find evidence for the cristalization phenomenon when increasing the genome-wide deleterious mutation rate beyond the drift barrier. In the following, we consider a deleterious mutation rate $\mu$, that is the number of lethal equivalents introduced per generation de novo into the recessive genes of a gamete. As an indicator of the fitness of a population or their gene pool, mutation burden $B(t)$ and disease incidence $P(t)$ were studied for populations of constant size $K$ over a total time span of 100 000100000100\,000100 000 generations [11]. The proportion of diploid individuals in the population that are biallelic for such lethal equivalents in at least one gene is the incidence or prevalence rate $P(t)$. The number of pathogenic alleles in the entire gene pool is defined as the mutation burden $B(t)$. All haplotypes in the gene pool can also be assigned to a class indicating the number of genes that harbour pathogenic alleles. By that definition, $c_{0}$ is the least loaded class containing only the haplotype without any mutations, $c_{1}$ consists of all haplotypes with exactly one affected gene, and so on. All simulations were performed in the discrete Wright-Fisher Model or with adaptive dynamics. Both frameworks were shown to yield comparable results [17]. Due to maximal selection ($s=1$) and total recessivity ($h=0$), a pathogenic or deleterious allele can also be referred to as a lethal equivalent that prevents propagation if both haplotypes of a gene are affected.
@@ 3.1. Mutation burden beyond the drift barrier
Note that the prevalence is independent from the number of loci $N$ and equals the probability that a mutation appears on a gamete at birth. In Figure 3 A we see that in the case of no recombination ($r=0$) the simulations match these considerations. In the case of full recombination ($r=1$) however the allele frequency and hence also the haploid mutation burden is lower that what was expected by Nei [23]. However in the absence of recombination, these equilibria, while they may remain stable for several thousand generations, are fragile. For example, in Figure 1 at around generation 20k we observe a transition to a higher mutation burden and prevalence within a few generations. Further transitions follow over a larger time frame in a stochastic manner. On the level of haplotypes, these transitions are a consequence of the extinction of the least loaded classes, which has also been referred to as clicks of Muller‘s ratchet (Figure 1 B). After the extinction of the $c_{0}$ class the new least loaded class $c_{1}$ of gametes with exactly one lethal equivalent is left without influx, but rather looses gametes due to de novo mutations. This would lead to a rapid fixation of one mutation within the population, even under weaker selection coefficients [5]. Fixation however is not possible without the extinction of the whole population for recessive lethals. Instead we observe that the extinction of the mutation free gamete sets off a cascade of extinction events that is only reassured by the formation of clusters of similar haplotypes that are mutually exclusive. This phenomenon of “cristalization” was already predicted for low dominance coefficients by Charlesworth and Charlesworth [5]. In the period after the population has stabilized, there can be instances of further genes being incorporated into one of the clusters. This may or may not be accompanied by the extinction of the least loaded class. Moreover, there can be a reduction in the number of clusters as they compete with each other. This results not only in an increase in mutation burden but also in a rapid rise in prevalence (see Figure 1 at around 50k).
@@ 4. Discussion
In this work we showed that the recessive gene count and recombination rate are important parameters in the drift barrier model that decide the stability of a gene pool. Beyond the barrier, the gene pool of a species becomes metastable, which means that transitions to higher levels of mutation burden can occur in a stochastic population model [2, 1]. The explanation on the level of haplotypes is the extinction of the least loaded classes that result in jumps of mutation burden and prevalence on much shorter time scales. Under real conditions, populations with such a high mutation burden would probably not be sustainable and die out.
@@
[…]
In our present work, we explored the interplay between different theories by examining how effective population size, genome-wide deleterious mutation rate, recessive gene count, and recombination rate might impact mutation burden and genome contamination. We did so by means of simulations with adaptive dynamics that behave equivalently to classical Wright-Fisher population genetics. We found the emergence of multiple mutually exclusive haplotypes, and for smaller numbers of genes, we often observed a “cristallization” into two complementary segregating haplotypes, a phenomenon predicted by Brian Charlesworth after studying earlier works from Pàlsson, et al. [3, 26, 25]. However, for larger numbers of recessive genes, we usually observed more than two clusters with a very strong heterozygote advantage at the end of our simulations representing 100 000100000100\,000100 000 generations. This represents most likely the fixed stage of Muller‘s ratchet after the repeated and irreversible loss of individuals containing the lowest number of deleterious mutations [5]. On the other hand, we extended the model of the drift barrier by showing that the recessive gene count has to be considered as an additional feature of the genomic architecture and that it is not enough to partition the genome into regions where deleterious mutations can occur. In an individual pathogenic variants in two different recessive genes are more harmful to the population than if they occurred in two different exons of the same gene, even though they contribute equally to the mutation burden. However, the complexity of the dynamics increases quickly with a growing gene count and we therefore focused on lethal equivalents in our simulations. In this case, the fitness function is simple, since it is a binary decision whether an individual is able to reproduce or not. In future work the effect of a wider range of selection coefficients and also more realistic recombination maps could be analysed.
@@ 7. Figures
Figure 1.

### v4

@@ 1. Introduction
While the drift-barrier model was originally based on empirical results, there has also been extensive theoretical work to understand the stochastic nature of this phenomenon. In simple terms, the dynamics of a gene pool are either stable or unstable, depending on which side of the barrier the system operates which is defined by its parameter settings (Figure 1). Most simulations in that field are based on the discrete Wright-Fisher Model or adaptive dynamics and both frameworks were shown to yield comparable results [21]. In our work, we consider a deleterious mutation rate $\mu$, that is the number of lethal equivalents introduced per generation de novo into the recessive genes of a gamete. As an indicator of the fitness of a population or their gene pool, mutation burden $B(t)$ and disease incidence $P(t)$ were studied for populations of constant size $K$ over a total time span of 100 000100000100\,000100 000 generations [15]. The proportion of diploid individuals in the population that are biallelic for such lethal equivalents in at least one gene is the incidence or prevalence rate $P(t)$. The number of pathogenic alleles in the entire gene pool is defined as the mutation burden $B(t)$. Incidence rate and mutation burden are also the system parameters that we use to assess its stability. All haplotypes in the gene pool can be assigned to a class indicating the number of genes that harbour pathogenic alleles. By that definition, $c_{0}$ is the least loaded class containing only the haplotype without any mutations, $c_{1}$ consists of all haplotypes with exactly one affected gene, and so on. Due to maximal selection ($s=1$) and total recessivity (dominance coefficient $h=0$), a pathogenic or deleterious allele can also be referred to as a lethal equivalent that prevents propagation if both haplotypes of a gene are affected. Muller first studied a simple stochastic process for lethal mutations in haploid genomes without recombination and observed the irreversible loss of the least loaded class from the population which he described as clicks of a ratchet [14, 26]. By theoretical arguments, Charlesworth and Charlesworth argued that in diploid genomes the accumulation of pathogenic or lethal recessive alleles can result in the “crystalization” of the population, that is, the occurrence of haplotypes that are incompatible with each other [5]. In our work, we study the stochastic process of Muller‘s ratchet in the limit of strong selection on completely recessive diploid genomes in a parameter space with a variable number of genes that operates close to the drift barrier. We find evidence for the crystalization phenomenon when increasing the genome-wide deleterious mutation rate or gene count beyond the drift barrier. A particular focus of our work are the inhomogeneous or metastable dynamics that follow after the extinction of the mutation-free gametes until the emergence of haploid clusters, when the population regains stability after several thousand generations [2]. Unlike in other models, we do not reach an equilibrium every time between two successive clicks [23]. In agreement with all other models, we can show that recombination weakens the selective disadvantage of a pathogenic recessive allele by exporting mutations to other members of the population [4, 7, 17, 19, 20]. Therefore, recombination is also the force that helps to balance genetic load to a certain degree in genomes with an increasing number of recessive genes, which we introduced as a novel parameter of the drift barrier.
@@ 3.1. Mutation burden beyond the drift barrier
Note that the prevalence is independent from the number of loci $N$ and equals the probability that a mutation appears on a gamete at birth. In Figure 4 A we see that in the case of no recombination ($r=0$) the simulations match these considerations. In the case of full recombination ($r=1$) however the allele frequency and hence also the haploid mutation burden is lower that what was expected by Nei [28]. However in the absence of recombination, these equilibria, while they may remain stable for several thousand generations, are fragile. For example, in Figure 2 at around generation 20k we observe a transition to a higher mutation burden and prevalence within a few generations. Further transitions follow over a larger time frame in a stochastic manner. On the level of haplotypes, these transitions are a consequence of the extinction of the least loaded classes, which has also been referred to as clicks of Muller‘s ratchet (Figure 2 B). After the extinction of the $c_{0}$ class the new least loaded class $c_{1}$ of gametes with exactly one lethal equivalent is left without influx, but rather looses gametes due to de novo mutations. This would lead to a rapid fixation of one mutation within the population, even under weaker selection coefficients [5]. Fixation however is not possible without the extinction of the whole population for recessive lethals. Instead we observe that the extinction of the mutation free gamete sets off a cascade of extinction events that is only reassured by the formation of clusters of similar haplotypes that are mutually exclusive. This phenomenon of “crystalization” was already predicted for low dominance coefficients by Charlesworth and Charlesworth [5]. In the period after the population has stabilized, there can be instances of further genes being incorporated into one of the clusters. This may or may not be accompanied by the extinction of the least loaded class. Moreover, there can be a reduction in the number of clusters as they compete with each other. This results not only in an increase in mutation burden but also in a rapid rise in prevalence (see Figure 2 at around 50k).
@@ 4. Discussion
In the present work, we aimed at approaching the drift barrier by numerical simulations for wide parameter ranges in all three dimensions. We did so by means of adaptive dynamics that behave equivalently to classical Wright-Fisher population genetics. We could also confirm the emergence of multiple mutually exclusive haplotypes, and for smaller numbers of genes. This “crystallization” into two complementary segregating haplotypes, is a phenomenon predicted by Brian Charlesworth after studying earlier works from Pàlsson, et al. [3, 31, 30]. In fact, we see our work as a new part of a long sequence of findings ever since John Haigh established a mathematical model in 1978 to quantify certain effects of Muller’s ratchet [14]. Particularly, the question of the click rate - that is, the speed at which successive least loaded classes become extinct - is of great interest. Recently, good approximations depending on effective population size have been achieved using diffusion approximation [9]. However, all these models observe time-homogeneous click rates. Interestingly, we observe highly inhomogeneous click rates and we hypothesized that this is due to the higher complexity of the genomic architecture that we modelled. This makes techniques like diffusion approximation or coalescent approaches from previous studies inapplicable [12, 1]. Indeed, after simplifying our model to only consider the number of deleterious mutations per gamete and assuming these mutations are always uniformly distributed across the genome in each step, we observed a homogeneous click rate again. Thus, the position of mutations is crucial, making mathematical analysis very challenging. The inhomogeneity is an interesting finding from a mere stochastic point of view and could be motivation enough to analyse the timespan between clicks and determine their distribution. Furthermore, for evolutionary biology it might be interesting to focus not only on the clicks of Muller’s ratchet but also on the addition of recessive genes to a cluster - which correlates with an increase in mutation burden - and the extinction of clusters - which correlates with an increase in prevalence. By this means our model can also contribute to the understanding of multilocus dynamics, that have recently been shown to the associative overdominance and background selection [11]. Providing these measures in terms of the order of effective population size would be a significant challenge. We share Charlesworth’s hypothesis that this evolution stops only when either the entire population becomes extinct or two exclusive clusters emerge, each encompassing the entire set of genes.
@@ 5. Code Availability
6. Further Engagement and Presentation
For those interested in a visually guided introduction to our main findings, I recently gave a short 15-minute talk at the European Society of Human Genetics (ESHG) conference in Berlin. This talk serves as an accessible entry point to our research, with the detailed context and insights provided in this paper. You can view the talk on YouTube via the following link: https://youtu.be/QXoHY5ODkVk
@@ 8. Figures
Figure 1. A shchematic representation of the Drift-Barrier in a three-dimensional parameter space. Below the hyperplane, populations can exist stably, whereas above it, the risk of extinction becomes too high. While the effect of the population size has already been described we observe an exponentially decreasing effect of the recessive gene count (see Figure 3). The combined effect has not yet been investigated. This is merely a sketch to visualize the extension of the drift-barrier hypothesis.